OpenStudy (anonymous):

Which of the following equations represents a perpendicular line to 6x - 2y = 20? y = 3x + 2 y = -3x + 4 y=1/3x-3 y=-1/3x-5

OpenStudy (anonymous):

So do you want to learn how to do it or just have an answer?

OpenStudy (anonymous):

how to do it

OpenStudy (anonymous):

first solve the original line for y :)

OpenStudy (anonymous):

6x - 2y = 20 sfy y = 3 x-10 I assume you know how to do this

OpenStudy (anonymous):

y = 3 x-10 since it is now in y= m x+b from and m is the slope which means 3 is the slope. another way of saying 3 is 3/1 =3 right... if you are given a slope then a slope or line perpendicular to is has to be flipped and multiplied by 1- so 3/1 become -1/3

OpenStudy (anonymous):

D y=-1/3x-5 is the only one with an m such as this ^m= -1/3

OpenStudy (anonymous):

just remember given any line and it's slope=rise/run then a slope perpendicular to it is -run/rise in your case 3 which is 3/1 became -1/3

OpenStudy (anonymous):

aalso note these drawings and perpindicular lines   THe point all of them are perpendicular and can be shifted left and right and still be perpendicular. In the future i think you will be given more points so that you have to find a specific line.